OpenStudy (anonymous):

x(x+2)= 24+4x i neeeeed helppp pleaseeee

OpenStudy (anonymous):

x^2+2x=24+4x x^2-2x-24=0 (x-6)(x+4)=0 x=6 or x=-4

OpenStudy (anonymous):

thank you very much